Do It! Review 25-3 The standard cost of product 5252 includes 1.90 hours of direct labor at $12.30 per hour. The predetermined overhead rate is $22.00 per direct labor hour. During July, the company incurred 4,000 hours of direct labor at an average rate of $12.60 per hour and $80,400 of manufacturing overhead costs. It produced 2,000 units. Compute the total, price, and quantity variances for labor. Total labor variance Labor price varñiance Labor quantity variance s Compute the total overhead variance Total overhead variance LINK TO TEXT By accessing this Question Assistance, you will learn while you eam points based on the Point Potential Policy set by your instructor. Question Attempts: 0 of 3 used SAVE POR LATER ! SUBHIT ANSWE
Answer #1

SOLUTION

A. Total labor variance = Total standard cost - Total actual cost

= (2,000*1.90 hours * $12.30) - (4,000*$12.60)

= $46,740 - $50,400

= $3,660 Unfavorable

Labor price variance = (Standard cost per hour - Actual hour per hour) * Actual hours

= ($12.30 - $12.60) * 4,000

= 0.30 * 4,000

= $1,200 Unfavorable

Labor quantity variance = (Standard hours - Actual hours) * Standard rate per hour

= ((2,000*1.90 hours) - 4,000) * $12.30

=(3,800 hours - 4,000 hours) * $12.30

= 200 hours * $12.30

= $2,460 Unfavorable

B. Total Overhead Variance= (Standard hours * Predetermined overhead rate) - Manufacturing overhead costs

= (2,000*1.90 hours * $22) - $80,400

= $83,600 - $80,400

= $3,200 Favorable